OpenStudy (anonymous):

Differential Eq: Find the general solution of the following ordinary differential equations (your answer will involve an unknown constant): x^2*y' = x*y + x*e^(y/x)

OpenStudy (anonymous):

\[x^2*y' = x*y + x*e ^{\frac{ y }{ x }}\]

OpenStudy (anonymous):

I'm kind of lost on how to approach this.

OpenStudy (kainui):

What have you tried?

OpenStudy (anonymous):

I tried separating it first, but I can't or don't know how I would cancel out e without getting y' stuck in the ln, and if I did that I would to apply e to both sides getting y/x stuck again.

OpenStudy (anonymous):

Just verified that its homogenous.

OpenStudy (anonymous):

In cases where you see something horrifying like \(e^{y(x) / x} \) it's often a good idea to define a new variable \( u(x) = y(x)/x\) and see how that goes.

OpenStudy (anonymous):

have u tried any substituion of variables

OpenStudy (anonymous):

I'm guessing I would have to set y=vx?

OpenStudy (anonymous):

say u = y/x

OpenStudy (anonymous):

y' = v+xv y'\[y' = \frac{ xy + x^2e^y/x }{ x^2 }\] \[v +xv' =\frac{ vx^2 +x^2e^v }{ x^2 } = v+e^v\] \[xv' = e^v\] \[\frac{ dv }{ dx } = \frac{ e^v }{ x }\]

OpenStudy (anonymous):

Then just separate and integrate Awesome. Thanks, err body

OpenStudy (anonymous):

Y'all are the real MVPs.

OpenStudy (anonymous):

That's it xD

ganeshie8 (ganeshie8):

\[\large y' = \frac{ xy + x^{\color{Red}{2}}e^{y/x} }{ x^2 }\] double check that exponent, your original equation has just \(x\)

OpenStudy (anonymous):

Oh damn. I missed that exponent in the original equation

ganeshie8 (ganeshie8):

either ways you will end up wid a separable eq'n, so not to wry too much.. just fix that mistake..

OpenStudy (anonymous):

thanks for looking out though.

OpenStudy (kainui):

For fun: Is it possible to solve this with Lambert's W function? =)

OpenStudy (dan815):

who does math for fun! eww nurd

OpenStudy (anonymous):

Lambert W function says that x = W(x)*exp(W(x)) right?

OpenStudy (kainui):

Yeah that's right.

OpenStudy (anonymous):

then we have 1 = W'(x)*exp(W(x)) + W(x)*W'(x)*exp(W(x)) = (1+W(x))*W'(x)*exp(W(x))

OpenStudy (anonymous):

and we have: v' = exp(v)/x, so we could say that v = W(x), then: \[\frac{1}{(1+W(x))e^{W(x)}} = \frac{e^{W(x)}}{W(x)e^{W(x)}}\] \[W(x) = e^{W(x)} + x\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

\[W^2(x) = x + xW(x)\] \[W(x) = \frac{x \pm \sqrt{x^2 + 4x}}{2}\]

OpenStudy (anonymous):

Mas W(x) = v(x) = y(x)/x, so: \[y(x) = \frac{x^2 \pm \sqrt{x^4 + 4x^2}}{2}\]